OpenStudy (anonymous):

How can I prove that for all symmetric matrices, its inverse is also symmetric?

myininaya (myininaya):

so we know \[A=A^T\]

myininaya (myininaya):

one sec

myininaya (myininaya):

ok lets me think

OpenStudy (anonymous):

yea we know \[A^{T}=A\] if symmetric. no problem, take your time. thanks.

myininaya (myininaya):

god joe is awesome at this stuff

OpenStudy (anonymous):

lol, i thought i had it, one sec <.<

myininaya (myininaya):

i need to go get my linear algebra book

OpenStudy (anonymous):

my linear algebra book doesn't have this proof. :(

myininaya (myininaya):

no zarkon is here we won't be able to do it now because he will i need to think faster

OpenStudy (zarkon):

\[(A^T)^{-1}=(A^{-1})^T\]

OpenStudy (zarkon):

add one step to what i have above and you have a proof

myininaya (myininaya):

\[A^T(A^{-1})^T=(A^{-1}A)^T=I^T=I\]

OpenStudy (zarkon):

that proves my statement

OpenStudy (anonymous):

But this would only work if A is orthonormal isn't it?

myininaya (myininaya):

\[(A^{-1})^TA^T=(AA^{-1})^T=I^T=I\]

OpenStudy (anonymous):

that are only using the fact that A is symmetrical. It doesnt have to be orthonormal.

myininaya (myininaya):

A has to be a n by n invertible matrix for the above to hold

myininaya (myininaya):

it says i spelled invertible wrong

myininaya (myininaya):

what am i missing

OpenStudy (zarkon):

\[A^{-1}=(A^T)^{-1}=(A^{-1})^T\]

myininaya (myininaya):

invertible i think tthats right

OpenStudy (anonymous):

give me a sec...i'm trying to understand it...haha...

OpenStudy (anonymous):

but i still don't understand how does this show that the inverse of a symmetric matrix will also be symmetric?

OpenStudy (zarkon):

I just showed that the inverse matrix is symmetric...it is equal to it transpose

OpenStudy (anonymous):

i would do it like this: Lets call the inverse of A B instead of A inverse (that -1 floating around everywhere is pissing me off). Then we have: \[AB = BA = I\] \[AB= I \iff (AB)^T = I^T \iff B^TA^T = I \iff B^TA = I\] so we have: \[B^TA =I , BA = I\] \[B^TA = BA \iff B^T = B\] Thus B is symmetric.

OpenStudy (zarkon):

that works too...just longer

OpenStudy (anonymous):

yeah, yours is good. I just hate that -1 floating around. And I also wanted to do this for myself lolol

OpenStudy (anonymous):

@zarkon, for the one you written: \[A^{-1}=(A^{T})^{-1}=(A^{-1})^{T}\] How can do I know from start that \[A^{-1}\] is \[A^{T}\] from the start?

OpenStudy (anonymous):

As in how is \[A^{-1}=(A^{T})^{-1}\]. I am not sure about this part.

OpenStudy (anonymous):

@joemath, for the part \[B^TA =I , BA = I\], why is \[B^{T}=B\]?

myininaya (myininaya):

B is symmetrical

OpenStudy (zarkon):

\[A^{-1}=(A)^{-1}\] \[A=A^T\] thus \[A^{-1}=(A)^{-1}=(A^T)^{-1}\] I'm just replacing A by A^T since they are equal

OpenStudy (anonymous):

oh ya you are so right! didn't thought of this! thanks for enlightening me!

OpenStudy (zarkon):

for what Joe wrote \[B^TA =I , BA = I\] this tells us that \[B^TA=BA\] multiply by \[A^{-1}\] on the right \[B^TAA^{-1}=BAA^{-1}\] thus \[B^T=B\]

OpenStudy (anonymous):

thanks for covering for me zarkon :)

OpenStudy (anonymous):

ohh now i got what Joe wrote! that's equally brilliant too. thanks guys!! really appreciate your help.